# CRISP-NAM: Competing Risks Interpretable Survival Prediction with Neural Additive Models

### v3

@@ 3.8 Interpretability Mechanisms
Given that CRISP-NAM is based on NAMs, as with all variants of Generalized Additive Models, we can use shape functions plots to visualize the (non-linear) contribution of each feature to the prediction. Specifically, for each feature $i$ and risk $k$, we can extract a shape function that describes how the feature affects the log-hazard ratio.
@@
Additionally, the importance of feature $i$ for risk $k$ is quantified by the mean absolute value of its contribution across the dataset.
@@ 4 Experiments
In this section, the datasets used and the experimental procedure to evaluate the CRISP-NAM model is described.
@@ 4.1 Datasets
In order to evaluate the proposed interpretable model for competing risks survival prediction, we used the following three real-world medical datasets:
@@
The PBC dataset originates from a randomized controlled trial conducted at the Mayo Clinic between 1974 and 1984, involving 312 patients diagnosed with primary biliary cholangitis. The study aimed to evaluate the efficacy of D-penicillamine in treating the disease. Each patient record includes 25 covariates encompassing demographic, clinical, and laboratory measurements. The primary endpoint was mortality while on the transplant waiting list, with liver transplantation considered a competing risk (Therneau and Grambsch, 2000).
@@
The SUPPORT2 dataset originates from the Study to Understand Prognoses and Preferences for Outcomes and Risks of Treatments (SUPPORT2), a comprehensive investigation conducted across five U.S. medical centers between 1989 and 1994. This dataset encompasses records of 9,105 critically ill hospitalized adults, each characterized by 42 variables detailing demographic information, physiological measurements, and disease severity indicators. The study was executed in two phases: Phase I (1989–1991) was a prospective observational study aimed at assessing the care and decision-making processes for seriously ill patients, while Phase II (1992–1994) implemented an intervention to enhance end-of-life care. The primary objective was to develop and validate prognostic models estimating 2- and 6-month survival probabilities, thereby facilitating improved clinical decision-making and patient-physician communication regarding treatment preferences and outcomes (Investigators et al., 1995).
@@
The dataset comprises 30,000 rows of unique patient data with 50% random right-censoring by drawing a censoring time $t_{c}^{(i)}\sim\mathcal{U}[0,\min\{T_{1}^{(i)},T_{2}^{(i)}\}]$. The final observed data for each patient is $(\mathbf{x}^{(i)},t^{(i)},k^{(i)})$, where $t^{(i)}$ is the observed time and $k^{(i)}$ is the event indicator ($\emptyset$ if censored).
@@ 4.2 Experimental Setup
The CRISP-NAM model is implemented using PyTorch. The experimental procedure is comprised of hyperparameter optimization followed by cross-validation evaluation phases. First, Optuna (Akiba et al., 2019) is employed to systematically search for optimal model configurations, tuning learning rate, $L_{2}$ regularization strength, dropout rates, network architecture (1-3 hidden layers with 32-256 units), and batch normalization settings using validation loss as the objective. The best configuration was then evaluated through stratified 5-fold cross-validation on multiple datasets previously described. Continuous features were normalized using standard scaling ($\mu=0,\sigma=1$). Missing values for categorical variables were imputed using mode imputation, and for continuous variables, mean imputation is applied for all datasets, except for the SUPPORT2 dataset for which median imputation was applied. Training employed the AdamW (Loshchilov and Hutter, 2017) optimizer to minimize the negative log-likelihood loss with a batch size of 256 and early stopping (patience=10) to prevent over-fitting.
@@ 5 Results and Discussion
CRISP-NAM achieves discrimination metrics within 2-5% of DeepHit on clinical datasets, demonstrating its competitiveness. On FHS, the TD-AUC gap at $q_{0.25}$ is merely 0.004 (0.850 vs 0.854), with CRISP-NAM actually outperforming DeepHit at $q_{0.50}$ for Risk 1 (0.834 vs 0.831). Similarly, on SUPPORT2 Risk 1, CRISP-NAM exceeds DeepHit’s TD-AUC by substantial margins at $q_{0.25}$ (0.894 vs 0.779) and $q_{0.50}$ (0.839 vs 0.640). For PBC, while DeepHit achieves marginally higher TD-AUC, both models reach near-ceiling performance ($\geq 0.96$), making the practical difference negligible. The synthetic dataset represents the primary challenge, where DeepHit’s flexible architecture captures complex non-linear patterns that CRISP-NAM’s additive structure cannot fully represent. While DeepHit generally achieves lower Brier scores, CRISP-NAM maintains clinically acceptable calibration on most datasets. The calibration gap is smallest on FHS (within 0.01-0.02) and moderate on PBC Risk 1. The larger gaps on SUPPORT2 and PBC Risk 2 reflect the trade-off between interpretability and flexibility in probability estimation. CRISP-NAM exhibits lower variance across folds in several scenarios (e.g., FHS Risk 1: ±0.009 vs ±0.012 for DeepHit at $q_{0.50}$), suggesting more stable predictions.
The performance gaps between CRISP-NAM and DeepHit are often clinically insignificant. A TD-CI difference of 0.02-0.03 rarely changes risk stratification decisions in practice. CRISP-NAM’s competitive performance, combined with its interpretability, positions it as a valuable alternative when model transparency is critical, particularly in high-stakes medical applications requiring explainable predictions.
@@ 5.1 Interpretability Analysis
We present shape function plots for the top 10 features per risk as learned by the CRISP-NAM model across 3 datasets: SUPPORT2, Framingham and PBC. Each curve $s_{i,k}(x_{i})$ represents the marginal contribution of feature $x_{i}$ to the log cause-specific hazard for risk $k$. Rug plots beneath each curve illustrate the empirical distribution of feature values, highlighting regions, particularly in the tails, where data are sparse.
Interpretation Note. These shape functions are associational, not causal, and may obscure interactions or time-varying effects. They are estimated under smoothness constraints and can extrapolate in regions with low data density, leading to amplified or flattened effects. As such, the shapes should be used primarily for model validation and debugging. Apparent patterns should be interpreted cautiously, corroborated on external cohorts, and discussed with domain experts before drawing scientific or clinical conclusions. While our primary focus centres on analyzing trends revealed through the shape function plots, we provide limited discussion of the associations between covariates and predicted risks. These discussions serve primarily to demonstrate how our shape plot findings align with or contrast against established medical literature.
@@ 5.1.1 Framingham Heart Study Dataset
Figure 2 presents shape functions from the CRISP-NAM model trained on the Framingham dataset, with separate plots for two competing risks: cardiovascular disease (Risk 1) and non–cardiovascular death (Risk 2). All functions are displayed on the log-hazard scale, where a unit increase of $+0.69$ corresponds to a doubling of the cause-specific hazard.
For Risk 1, AGE, SYSBP, BMI, GLUCOSE, and the binary indicator DIABETES exhibit steadily increasing trends across most of their observed ranges. DIABP and AGE display similar steep upward slopes beyond their respective median values. TOTCHOL shows a mild U-shaped profile with a slight increase at high values. educ has an approximately linear decreasing trend across its range. The binary history flags PREVCHD and PREVAP show negative contributions to the log-hazard.Three data characteristics, rather than a real reversal of risk, explain this result:
@@
Several observed patterns in the shape functions are consistent with known risk factors for cardiovascular and all-cause mortality. Established studies have linked increasing age, elevated systolic blood pressure, high glucose levels, and diabetes with heightened cardiovascular risk (Kannel and McGee, 1979; D’Agostino and Kannel, 1989). The inverse trend for educational attainment aligns with literature on socioeconomic disparities in cardiovascular outcomes (Rosengren et al., 2019). The U-shaped relationship observed for BMI (Risk 2) has been previously noted in older adults and is often described as the “obesity paradox” (Amundson et al., 2010) as well as sparse data in the extreme ranges. Similarly, the positive association between elevated diastolic pressure and non-cardiovascular mortality may reflect constraints due to cohort truncation and sparsity in the extreme blood-pressure ranges.
@@ 5.1.2 Primary Biliary Cholangitis (PBC) Dataset
For Risk 1, age exhibits a monotonic increasing contribution to the predicted hazard. SGOT and alkaline display sharp increases followed by plateaus at higher values. The function for serBilir rises steeply at low values and saturates thereafter. serChol peaks at moderate levels and decreases subsequently. Binary indicators such as hepatomegaly_Yes and edema_edema despite diuretics show positive contributions. albumin demonstrates an inverted S-shaped curve: initially increasing, crossing zero near the median, and declining at higher values. platelet decreases steadily, with a minor upturn in the extreme right tail where data are sparse.
For Risk 2, the shape for age decreases monotonically. SGOT rises rapidly and then flattens. histologic shows an approximately linear increasing trend. The shape for serBilir forms an inverted-U. albumin rises beyond the cohort median and levels off at higher values. ascites_Yes and hepatomegaly_Yes are associated with negative contributions. alkaline increases sharply before flattening. spiders_Yes exhibits a slight negative effect. Finally, serChol increases with higher values, especially in the upper tail, although data in that region are sparse.
The rug plots accompanying each shape function reflect the distribution of the feature values and indicate regions with limited data support.These empirical patterns align with several well-established clinical insights in the context of Primary Biliary Cholangitis (PBC). For instance, older age, elevated liver enzymes such as SGOT and alkaline phosphatase, and increased serum bilirubin are recognized markers of disease severity and poorer prognosis (Dickson et al., 1989; Murillo Perez et al., 2023). The decreasing transplant hazard with age may reflect clinical prioritization criteria that favour younger candidates for organ allocation. The non-linear shape for albumin aligns with its known role as a proxy for liver synthetic function, where low levels indicate hepatic decompensation. Histologic stage progression, from fibrosis to cirrhosis, is a standard determinant in transplant eligibility, consistent with the monotonic rise observed in its shape function. Additionally, ascites and hepatomegaly are classical signs of advanced liver disease, often associated with higher mortality and reduced transplant suitability. The distinct inverted-U pattern for bilirubin in the transplant hazard may reflect clinical decision-making thresholds, while the rising contribution of cholesterol could be interpreted in light of its association with bile duct dysfunction in PBC.
@@ 5.1.3 SUPPORT2 Dataset
Figure 6 presents feature importance and shape functions from the CRISP-NAM model trained on the SUPPORT2 dataset, which distinguishes cancer-specific mortality (Risk 1) from death due to other causes (Risk 2). All shape functions are plotted on the log-hazard scale.
For Risk 1, the variable sps shows a near-linear increase, indicating a progressive rise in the predicted hazard across its range. The shape function for age follows an inverted-U profile, with risk increasing up to approximately 65 years and declining thereafter. Binary indicators dementia_1 and dnr_no_dnr exhibit negative contributions to the log-hazard. The shape for avtisst is flat at lower durations of mechanical ventilation but shows a steep increase beyond approximately 3 days. Negative contributions for dnr_no_dnr and dementia_1 may reflect treatment-limiting decisions or documentation timing effects, as suggested in earlier SUPPORT studies (Hakim et al., 1996).
For Risk 2, the severity scores aps and sps both show sharp increases followed by plateaus. The avtisst variable again displays a marked rise in hazard for durations exceeding 3 days. Cost-related variables charges and totcst show decreasing trends. The binary indicator dnr_no_dnr also shows a downward slope across its two levels. Rising sps and aps scores are consistent with the role of physiological instability and organ dysfunction in predicting mortality (Knaus et al., 1995). The inverted-U for age in cancer mortality likely reflects competing risks from other causes in older individuals (Hayes-Larson et al., 2020). The association between prolonged ventilation (avtisst $>3$ days) and increased hazard aligns with the known severity of illness in patients requiring extended respiratory support. Cost variables likely act as proxies for length of stay or illness trajectory rather than direct predictors.

### v5

@@ 3.8 Interpretability Mechanisms
Given that CRISP-NAM is based on NAMs, as with all variants of Generalized Additive Models, CRISP-NAM can generate shape functions plots to visualize the (non-linear) contribution of each feature to the prediction. Specifically, for each feature $i$ and risk $k$, we can extract a shape function that describes how the feature affects the log-hazard ratio.
@@
The importance of feature $i$ for risk $k$ is quantified by the mean absolute value of its contribution across the dataset.
@@
Notably, in this current implementation, CRISP-NAM does not capture features interactions. This is by design to prioritize interpretability through independent feature level shape functions, ensuring that feature contributions can be visualized and understood in isolation. Adding separate FeatureNets to model feature interactions adds to the model complexity and affects its interpretability as visualization beyond pairwise features interactions is challenging. With $p$ features, there are $\mathcal{O}(p^{2})$ possible pairwise interactions, and deciding which interactions to model would also require domain knowledge.
@@ 4 Experiments
In this section, the datasets used and the experimental procedure to evaluate the CRISP-NAM model are described.
@@ 4.1 Datasets
In order to evaluate the proposed interpretable model for competing risks survival prediction, we used the following three real-world medical datasets and a synthetic dataset. Table 1 provides a summary and breakdown of the primary and competing risks for each the datasets used in this work.
@@
The PBC dataset originates from a randomized controlled trial conducted at the Mayo Clinic between 1974 and 1984, involving 312 patients diagnosed with primary biliary cholangitis. The study aimed to evaluate the efficacy of D-penicillamine in treating the disease. Each patient record includes 25 covariates encompassing demographic, clinical, and laboratory measurements. The primary endpoint was mortality while on the transplant waiting list, with liver transplantation considered a competing risk (Dickson et al., 1989).
@@
The SUPPORT2 dataset originates from the Study to Understand Prognoses and Preferences for Outcomes and Risks of Treatments (SUPPORT2), a comprehensive investigation conducted across five U.S. medical centers between 1989 and 1994. This dataset encompasses records of 9,105 critically ill hospitalized adults, each characterized by 42 variables detailing demographic information, physiological measurements, and disease severity indicators. The study was executed in two phases: Phase I (1989–1991) was a prospective observational study aimed at assessing the care and decision-making processes for seriously ill patients and Phase II (1992–1994) implemented an intervention to enhance end-of-life care. The primary objective was to develop and validate prognostic models estimating 2- and 6-month survival probabilities, thereby facilitating improved clinical decision-making and patient-physician communication regarding treatment preferences and outcomes (sup, 1995).
@@
The dataset consists of 30,000 rows of unique patient data with 50% random right-censoring by drawing a censoring time $t_{c}^{(i)}\sim\mathcal{U}[0,\min\{T_{1}^{(i)},T_{2}^{(i)}\}]$. The final observed data for each patient is $(\mathbf{x}^{(i)},t^{(i)},k^{(i)})$, where $t^{(i)}$ is the observed time and $k^{(i)}$ is the event indicator ($\emptyset$ if censored).
Table 1: Dataset characteristics and competing risk statistics
Dataset Observations Features Primary Competing Risk Censored
PBC 312 25 Death (44.87%) Transplant (9.29%) 45.83%
Framingham 4,434 18 CVD (26.09%) Death (17.75%) 56.16%
SUPPORT2 9,105 42 Cancer_Death (18.2%) Death_Other (49.9%) 31.9%
Synthetic 30,000 12 * (25.33%) * (24.67%) 50.00%
@@ 4.2 Experimental Setup
The CRISP-NAM model is implemented using PyTorch and the code is available from https://github.com/VectorInstitute/crisp-nam. We employed nested cross-validation to prevent data leakage during hyperparameter optimization and model evaluation. The approach consists of an outer 5-fold stratified cross-validation for performance evaluation and an inner 5-fold cross-validation for hyperparameter tuning within each outer fold. For each outer fold, Optuna (Akiba et al., 2019) is employed on the training partition to systematically search for optimal model configurations using the inner 5-fold cross-validation, tuning learning rate, $L_{2}$ regularization strength, dropout rates, network architecture (1-3 hidden layers with 8-128 units), and batch normalization settings using validation loss as the objective. The best configuration identified for each outer fold is then trained on the complete training partition and evaluated on the corresponding held-out test partition. Continuous features were normalized using standard scaling ($\mu=0,\sigma=1$) and categorical features were one-hot encoded. Missing categorical values were imputed using mode imputation. For continuous variables, mean imputation was used across all datasets. Training employed the AdamW (Loshchilov and Hutter, 2019) optimizer to minimize the negative log-likelihood loss with a batch size of 256 and early stopping (patience=10) to prevent over-fitting.
@@ 5 Results and Discussion
[…]
Notably, CRISP-NAM’s calibration performance (Brier score) is also lower than NFG and DeepHit. This could be attributed to the loss function we implemented that optimizes for discrimination (ranking) rather than the calibration (probability accuracy). In addition, the event weighting could exacerbate the model’s lagging calibration. We will investigate improvements to the loss function to handle calibration more effectively in our future work.
@@ 5.1 Interpretability Analysis
Here, the shape function plots for the top 10 features per risk as learned by the CRISP-NAM model across 3 datasets: SUPPORT2, Framingham and PBC are presented. Each curve $s_{i,k}(x_{i})$ represents the marginal contribution of feature $x_{i}$ to the log cause-specific hazard for risk $k$. Rug plots beneath each curve illustrate the empirical distribution of feature values, highlighting regions, particularly in the tails, where data are sparse.
Interpretation Note. It should be noted that these shape functions are associational, not causal, and may obscure interactions between features. They are estimated under smoothness constraints and can extrapolate in regions with low data density, leading to amplified or flattened effects. Apparent patterns should be interpreted cautiously, corroborated on external cohorts, and discussed with domain experts before drawing scientific or clinical conclusions. While our primary focus centres on analyzing trends revealed through the shape function plots, we provide limited discussion of the associations between covariates and predicted risks. These discussions serve primarily to demonstrate how our shape plot findings align with or contrast against established medical literature.
@@ 5.1.1 Framingham Heart Study Dataset
Figure 3 illustrates shape functions from the CRISP-NAM model trained on the Framingham dataset, with separate plots for two competing risks: cardiovascular disease (Risk 1) and non–cardiovascular death (Risk 2). All functions are displayed on the log-hazard scale, where a unit increase of +0.69+0.69 corresponds to a doubling of the cause-specific hazard.
For Risk 1, GLUCOSE and BMI shows steep increase showing that higher blood glucose levels and increasing BMI are associated to higher risk of cardiovascular disease for this dataset. Shape plots reveal that the patients with diabetes, patients who are on medication for hypertension and who are current smokers (DIABETES, BPMEDS and CURSMOKE) have higher risk of cardiovascular disease which correlates well with known risk factors. Shape plot also reveals that women have lower risk of cardiovascular disease, compared to men and having more education is correlated with lower risk of cardiovascular disease. The binary history flags PREVCHD and PREVAP show negative contributions to the log-hazard. Three data characteristics, rather than a real reversal of risk, explain this result:
@@
Several observed patterns in the shape functions are consistent with known risk factors for cardiovascular and all-cause mortality. Established studies have linked elevated systolic blood pressure and high glucose levels with heightened cardiovascular risk (Kannel and McGee, 1979; D’Agostino and Kannel, 1989). Cigarettes per day displayed an exponential relationship with diminishing marginal effects at higher consumption levels, reflecting saturation of smoking-related harm pathways. The inverse trend for educational attainment aligns with literature on socioeconomic disparities in cardiovascular outcomes (Rosengren et al., 2019). The U-shaped relationship observed for BMI (Risk 2) has been previously noted in older adults and is often described as the “obesity paradox” (Amundson et al., 2010) as well as sparse data in the extreme ranges. Additionally, previous cardiovascular conditions (angina pectoris and coronary heart disease) showed positive linear associations with mortality risk.
@@ 5.1.2 Primary Biliary Cholangitis (PBC) Dataset
For Risk 1, age exhibits a Sigmoid-shaped curve. alkaline displays sharp increases followed by plateaus at higher values.Biomarkers serBilir, serChol and prothrombin all show similar rises steeply at low values and plateus thereafter. Binary indicators such as hepatomegaly_Yes shows positive contributions. albumin demonstrates an inverted Sigmoid-shaped curve: initially increasing, crossing zero near the median, and declining at higher values.
For Risk 2, the shape for age exhibits inverse sigmoidal shape with a decreasing trend as age increases signifying that transplantation risk decreases as patients progress with age. The shape for serBilir, alkaline and serChol rises steeply, before plateauing. These three elevated biomarkers indicate disease severity in PBC, which simultaneously increases both death risk and transplant priority. Examining the contribution scales for both risks for ascites_Yes shows a moderate negative contribution to Risk 1 ($\sim{-0.4}$) but a much stronger negative contribution to Risk 2 ($\sim{-1.0}$). From a model validation perspective, this suggests the model has learned that ascites presence is associated with reduced likelihood of both outcomes, but particularly transplantation. The asymmetric magnitudes indicate the model distinguishes between the two competing risks rather than simply treating ascites as a general severity marker.
The rug plots accompanying each shape function reflect the distribution of the feature values and indicate regions with limited data support. These empirical patterns align with several well-established clinical insights in the context of Primary Biliary Cholangitis (PBC). For instance, older age, elevated liver enzymes such as alkaline phosphatase, and increased serum bilirubin are recognized markers of disease severity and poorer prognosis (Dickson et al., 1989; Murillo Perez et al., 2023). The decreasing transplant hazard with age may reflect clinical prioritization criteria that favour younger candidates for organ allocation. The non-linear shape for albumin aligns with its known role as a proxy for liver synthetic function, where low levels indicate hepatic decompensation. Histologic stage (histologic) progression, from fibrosis to cirrhosis, is a standard determinant in transplant eligibility, consistent with the monotonic rise observed in its shape function. Additionally, ascites and hepatomegaly are classical signs of advanced liver disease, often associated with higher mortality and reduced transplant suitability. Edema in PBC patients indicates advanced liver disease. In the competing risks framework, patients with edema have higher disease severity and thus receive higher priority for transplantation due to their urgent medical need. The positive contribution to transplantation risk reflects the clinical reality that transplant allocation systems prioritize sicker patients and those with edema are more likely to receive transplants because it serves as a marker of advanced disease requiring urgent intervention.
@@ 5.1.3 SUPPORT2 Dataset
Figure 7 presents feature importance and shape functions from the CRISP-NAM model trained on the SUPPORT2 dataset, which distinguishes cancer-specific mortality (Risk 1) from death due to other causes (Risk 2).
We highlight several observations from the shape plots for both risks. For Risk 1, the shape function for age shows a lower risk of death from cancer for younger patients, then steadily increasing risk of cancer-related death up to approximately 65 years and declining risk for very old patients. The binary indicator race_black exhibits a slight negative contribution to the log-hazard for Risk 1 showing that black patients appear to have lower mortality due to cancer. The apparent protective effect of Black race contradicts established epidemiological evidence (Saka et al., 2025) demonstrating higher cancer mortality rates in Black population. This could indicate insufficient sample representation or selection biases inherent to the SUPPORT2 dataset. Similarly, the inverse relationship between number of comorbidities and cancer death risk suggests competing mortality mechanisms, where patients with multiple comorbidities may succumb to other medical conditions before cancer progression becomes the primary threat. The shape for avtisst is lower durations of mechanical ventilation but shows a steep increase beyond certain number of days. pafi (oxygen ratio) exhibits a spike at very low values ($\sim{0.7}$), then drops to steady negative contribution ($\sim{0.3}$). Other biomarkers such as sodium sod shows an inverted U-shape peaking near normal levels, suggesting both low sodium levels as well as very high sodium levels increases risk.
For Risk 2, the severity scores aps show sharp increases followed by plateaus. The avtisst variable again displays a marked rise in hazard for durations exceeding 3 days. Cost-related variables charges and totcst show decreasing trends. The binary indicator dnr_no_dnr shows that patients without DNR (Do Not Resuscitate) orders (dnr_no_dnr = 1) demonstrated substantially lower contributions to non-cancer death risk compared to those with DNR orders present. This pattern aligns with clinical expectations, as DNR orders typically indicate patients with poor overall prognosis, advanced chronic diseases, or end-stage conditions who are at elevated risk for cardiovascular, respiratory, or multi-organ failure. Rising aps scores are consistent with the role of physiological instability and organ dysfunction in predicting mortality (Knaus et al., 1995). The inverted-U for age in cancer mortality likely reflects competing risks from other causes in older individuals (Hayes-Larson et al., 2020). The association between prolonged ventilation (avtisst >3>3 days) and increased hazard aligns with the known severity of illness in patients requiring extended respiratory support. Cost variables likely act as proxies for length of stay or illness trajectory rather than direct predictors.
@@ 6 Conclusion
We acknowledge that CRISP-NAM inherits the proportional hazards assumption from the Cox framework, requiring that covariate effects on each cause-specific hazard remain constant over time. This assumption may be violated when covariate effects vary temporally, such as biomarkers having different predictive power for early versus late events. There are two possible avenues for future work: (i) exploration of temporal FeatureNets capable of learning how feature contributions evolve over time. The caveat that is that this approach would likely increase computational complexity during training and potentially demand larger datasets. Additionally, we will investigate using a modified loss function that takes calibration (Brier Score) into account.